# Analysis of Decentralized Stochastic Successive Convex Approximation for composite non-convex problems

### v1

@@ IV Experimental data and results
where $\xi_{i}=n_{i}\sim\mathcal{N}(0,1)$. The objective function is non-convex, and its surrogate can be constructed using (2). First, D-MSSCA is used to demonstrate the effect of communication topology in Fig 1. We observe that the fully connected network performs better than the Tree network, which is how the behavior is expected. Next, setting $\alpha=0.8$ and $\beta=0.16$ and $\mu=5000$, we plot the evolution of local variable $x_{i}^{t}$ with different initial values in Fig 2 for a fully connected graph with ${\lambda_{{\mathbf{W}}}}=0.5$. We observe the nodes converge to local minima. Finally, we plot the evolution of local variables of each node given a global constraint set $\lvert x_{i}\rvert\leq 2.25$, when all the nodes are initialized at $x_{i}^{t}=0$ in Fig 3. It can be observed that all the local variables get as close as possible to the true minima.
@@ V Conclusion and Future work
In this work, we consider decentralized consensus stochastic non-convex optimization to minimize the sum of non-convex (possibly smooth) and convex (possibly non-smooth) cost functions over a network of nodes. While this problem is well studied, comprehensive convergence analysis under the stochastic SCA rubric has remained an open problem. We have analyzed and proposed D-MSSCA that achieves the optimal rate of $\mathcal{O}(\epsilon^{-3/2})$. The rate matches the SFO rate of the state-of-the-art decentralized gradient-based algorithm while processing a single sample at each iteration. The algorithm uses strong convex surrogates and leverages recursive momentum-based updates at each node, achieving faster convergence. The applicability of D-MSSCA is demonstrated on a synthetic stochastic problem. One interesting future direction of this paper is under investigation, wherein we simplify the optimization problem (1) in Algorithm 1 by linearizing the global constraint $g$ making it (1) easier to solve than proximal-based methods. By reducing the complexity of each iteration, this modification could improve convergence rates and expand the applicability of the D-MSSCA algorithm to broader classes of optimization problems, especially those where proximal methods face scalability and efficiency challenges.
@@ Appendix A Proof of Lemma 2
In (a), we have applied the property of ${\mathbf{W}}$, i.e., $\mathbf{1}^{\mathsf{T}}{\mathbf{W}}=\mathbf{1}^{\mathsf{T}}$. In (b), Young’s inequality and the property of norm are utilized, i.e., for any square matrix ${\mathbf{A}}$ and vector ${\mathbf{x}}$ of compatible size, $\left\|{\mathbf{A}}{\mathbf{x}}\right\|^{2}\leq\lambda_{\max}({\mathbf{A}})^{2}\left\|{\mathbf{x}}\right\|^{2}$. In (c), we have utilized (14a) and the definition of ${\lambda_{{\mathbf{W}}}}$.
@@ Appendix B Proof of Lemma 4
From the definition of $\varepsilon^{t}$ and (8)
@@
Now we will simplify each term in the above equation separately, starting with $\left\|{\mathbf{z}}^{t}-{\mathbf{z}}^{t-1}\right\|^{2}$ we have
@@
On adding and subtraction by $\beta\nabla u_{i}({\mathbf{x}}^{t-1})$ and simplifying further we get
@@
On taking expectation on both sides and further utilizing Assumptions A3 and A4 we get
${\mathbb{E}}\left[\left\|{\mathbf{z}}^{t}-{\mathbf{z}}^{t-1}\right\|^{2}\right]\leq 3L^{2}{\mathbb{E}}\left[\left\|{\mathbf{x}}^{t}-{\mathbf{x}}^{t-1}\right\|^{2}\right]+3\beta^{2}{\mathbb{E}}\left[\left\|\left({\mathbf{z}}^{t-1}-\nabla u({\mathbf{x}}^{t-1})\right)\right\|^{2}\right]+3\beta^{2}\bar{\sigma}^{2}.$ (28)
By using conditional expectation, we can further simplify the last term of (26) as
@@
By applying (27) and utilizing Assumption A2, we have ${\mathbb{E}}\left[{\mathbf{z}}^{t}-{\mathbf{z}}^{t-1}|{\mathcal{H}}_{k}\right]=\nabla u({\mathbf{x}}^{t})-\nabla u({\mathbf{x}}^{t})-\beta({\mathbf{z}}^{t-1}-\nabla u({\mathbf{x}}^{t-1}))$ Substituting this and applying the Cauchy-Schwarz inequality, we get
@@
On further applying Young’s inequality and a few mathematical simplifications, we get
@@ Appendix C Proof of Lemma 6
On substituting $\eta_{1}=\frac{1-{\lambda_{{\mathbf{W}}}}^{2}}{2{\lambda_{{\mathbf{W}}}}^{2}}$ in Lemma 2, we get
@@ Appendix E Proof of Lemma 8
On further combining the common terms, we get the desired result. ∎
@@ Appendix F Proof of Lemma 9
where $\hat{{\mathbf{w}}}_{i}^{t}\in\partial(h+11_{{\mathcal{X}}})\mid_{{\mathbf{x}}={\hat{{\mathbf{x}}}}_{i}^{t}}$, and $11_{{\mathcal{X}}}({\mathbf{x}})$ is an indicator function. $011_{{\mathcal{X}}}({\mathbf{x}})=$ 0 if ${\mathbf{x}}\in{\mathcal{X}}$, otherwise $11_{{\mathcal{X}}}({\mathbf{x}})=\infty$. As per the definition of ${\tilde{f}}\left({\mathbf{x}}_{i},{\mathbf{x}}_{i}^{t},\xi_{i}^{t}\right)$ given in (2) and using (4), we have $\nabla{\tilde{f}}({\mathbf{x}}_{i}^{t},{\mathbf{x}}_{i}^{t},{\boldsymbol{\xi}}_{i}^{t})={\mathbf{z}}_{i}^{t}$. Furthermore, by adding and subtracting $\nabla{\tilde{f}}({\mathbf{x}}_{i}^{t},{\mathbf{x}}_{i}^{t},{\boldsymbol{\xi}}_{i}^{t})={\mathbf{z}}_{i}^{t}$ and substituting $\pi_{i}^{t}={\mathbf{y}}_{i}^{t}-{\mathbf{z}}_{i}^{t}$, we get
@@
From the definition of ${\tilde{f}}\left({\mathbf{x}}_{i},{\mathbf{x}}_{i}^{t},\xi_{i}^{t}\right)$ (2), and utilizing (4), we have $\nabla{\tilde{f}}\left({\hat{\mathbf{x}}}_{i}^{t},{\mathbf{x}}_{i}^{t},\xi_{i}^{t}\right)-{\mathbf{z}}_{i}^{t}=\mu_{i}({\hat{{\mathbf{x}}}}_{i}^{t}-{\mathbf{x}}_{i}^{t})$, substituting this gives us:
@@
In $(i)$ and $(ii)$, we have utilized the convexity of $h+11_{\mathcal{X}}$, property of ${\mathbf{W}}$ being doubly stochastic, and $W_{i,j}>0$ for all $i,j\in{\mathcal{V}}$ ((A7)).
@@
On dividing (31) by $\alpha$ and adding in (30) we get
@@
$\leq\frac{1}{\alpha n}\sum_{i=1}^{n}\left(h\left({\mathbf{x}}_{i}^{t}\right)\right)+\frac{1}{n}\sum_{j=1}^{n}\langle\hat{{\mathbf{w}}}_{j}^{t},{\hat{{\mathbf{x}}}}_{i}^{t}-{\mathbf{x}}_{i}^{t}\rangle+\frac{-1}{n}\sum_{i=1}^{n}\mu\left\|{\hat{{\mathbf{x}}}}_{i}^{t}-{\mathbf{x}}_{i}^{t}\right\|^{2}.$
@@
On further applying Cauchy Schwartz inequality and peter-paul’s inequality for $\gamma_{1}>0$ we get,
@@
On taking the Expectation on both side and summing for all $1\leq t\leq T$
@@
Further from the zeroth order convexity condition of $h({\mathbf{x}})$ we have $h({\bar{{\mathbf{x}}}}^{T+1})\leq\frac{1}{n}\sum_{i=1}^{n}h({\mathbf{x}}_{i}^{T+1})$ and from the initialization of D-MSSCA we have ${\mathbf{x}}_{i}^{1}={\bar{{\mathbf{x}}}}^{1}$ for all $i\in{\mathcal{V}}$, utilizing which we get the desired result.

### v2

@@ IV Experimental data and results
where $\xi_{i}=n_{i}\sim\mathcal{N}(0,1)$. The objective $U(x)$ in the interval $[-4,4]$ is shown in Fig. 1. As the objective function is non-convex, and its surrogate can be constructed using (2). We use D-MSSCA to first demonstrate the effect of communication topology in Fig 2. We observe that the fully connected network performs better than the Tree network, which is how the behavior is expected. Next, setting $\alpha=0.8$ and $\beta=0.16$ and $\mu=5000$, we plot the evolution of local variable $x_{i}^{t}$ with different initial values in Fig 3 for a fully connected graph with ${\lambda_{{\mathbf{W}}}}=0.5$. We observe the nodes converge to local minima. Finally, we plot the evolution of local variables of each node given a global constraint set $\lvert x_{i}\rvert\leq 2.25$, when all the nodes are initialized at $x_{i}^{t}=0$ in Fig 4. It can be observed that all the local variables get as close as possible to the true minima.
Figure 1: Golbal objective function $U(x)$.
@@ V Conclusion and Future work
In this work, we consider decentralized consensus stochastic non-convex optimization to minimize the sum of non-convex (possibly smooth) and convex (possibly non-smooth) cost functions over a network of nodes. While this problem is well studied, comprehensive convergence analysis under the stochastic SCA rubric has remained an open problem. We have proposed and analyzed D-MSSCA that achieves the optimal rate of $\mathcal{O}(\epsilon^{-3/2})$. The rate matches the SFO rate of the state-of-the-art decentralized gradient-based algorithm while processing a single sample at each iteration. The algorithm uses strong convex surrogates and leverages recursive momentum-based updates at each node, achieving faster convergence. The applicability of D-MSSCA is demonstrated on a synthetic stochastic problem. One interesting future direction of this paper is under investigation, wherein we simplify the optimization problem (1) in Algorithm 1 by linearizing the global constraint $g$ making it (1) easier to solve than proximal-based methods. By reducing the complexity of each iteration, this modification could improve convergence rates and expand the applicability of the D-MSSCA algorithm to broader classes of optimization problems, especially those where proximal methods face scalability and efficiency challenges.
@@ Appendix A Proof of Lemma 2
In (a), we have applied the property of ${\mathbf{W}}$, i.e., $\mathbf{1}^{\mathsf{T}}{\mathbf{W}}=\mathbf{1}^{\mathsf{T}}$. In (b), Young’s inequality and the property of norm are used, i.e., for any square matrix ${\mathbf{A}}$ and vector ${\mathbf{x}}$ of compatible size, $\left\|{\mathbf{A}}{\mathbf{x}}\right\|^{2}\leq\lambda_{\max}({\mathbf{A}})^{2}\left\|{\mathbf{x}}\right\|^{2}$. In (c), we have applied (14a) and the definition of ${\lambda_{{\mathbf{W}}}}$. ∎
@@ Appendix B Proof of Lemma 4
From the definition of $\varepsilon^{t}$ and (8) we have,
@@
Where (i), uses the doubly stochastic matrix property of ${\mathbf{W}}$. Next, each term in the above equation will be simplified separately, starting with $\left\|{\mathbf{z}}^{t}-{\mathbf{z}}^{t-1}\right\|^{2}$ we have,
@@
After adding and subtracting $\beta\nabla u_{i}({\mathbf{x}}^{t-1})$, we observe that
@@
Taking expectation on both sides and further using Assumptions A3 and A4 we get
${\mathbb{E}}\left[\left\|{\mathbf{z}}^{t}-{\mathbf{z}}^{t-1}\right\|^{2}\right]\leq 3L^{2}{\mathbb{E}}\left[\left\|{\mathbf{x}}^{t}-{\mathbf{x}}^{t-1}\right\|^{2}\right]+3\beta^{2}{\mathbb{E}}\left[\left\|{\mathbf{z}}^{t-1}-\nabla u({\mathbf{x}}^{t-1})\right\|^{2}\right]+3\beta^{2}\bar{\sigma}^{2}.$ (28)
Considering conditional expectation, we can further simplify the last term of (26) as
@@
Moreover, using (27) and Assumption A2, we have ${\mathbb{E}}\left[{\mathbf{z}}^{t}-{\mathbf{z}}^{t-1}|{\mathcal{H}}_{k}\right]=\nabla u({\mathbf{x}}^{t})-\nabla u({\mathbf{x}}^{t})-\beta({\mathbf{z}}^{t-1}-\nabla u({\mathbf{x}}^{t-1}))$. Substituting which along with Cauchy Schwartz inequality, gives
@@
Young’s inequality along with few mathematical simplifications gives,
@@ Appendix C Proof of Lemma 6
Substituting $\eta_{1}=\frac{1-{\lambda_{{\mathbf{W}}}}^{2}}{2{\lambda_{{\mathbf{W}}}}^{2}}$ in Lemma 2, we get
@@ Appendix E Proof of Lemma 8
Further combining the common terms, we get the desired result. ∎
@@ Appendix F Proof of Lemma 9
where $\hat{{\mathbf{w}}}_{i}^{t}\in\partial(h+11_{{\mathcal{X}}})\mid_{{\mathbf{x}}={\hat{{\mathbf{x}}}}_{i}^{t}}$, and $11_{{\mathcal{X}}}({\mathbf{x}})$ is an indicator function. As per the definition of ${\tilde{f}}\left({\mathbf{x}}_{i},{\mathbf{x}}_{i}^{t},\xi_{i}^{t}\right)$ given in (2) and using (4), we have $\nabla{\tilde{f}}({\mathbf{x}}_{i}^{t},{\mathbf{x}}_{i}^{t},{\boldsymbol{\xi}}_{i}^{t})={\mathbf{z}}_{i}^{t}$. Furthermore, by adding and subtracting $\nabla{\tilde{f}}({\mathbf{x}}_{i}^{t},{\mathbf{x}}_{i}^{t},{\boldsymbol{\xi}}_{i}^{t})={\mathbf{z}}_{i}^{t}$ and substituting $\pi_{i}^{t}={\mathbf{y}}_{i}^{t}-{\mathbf{z}}_{i}^{t}$, we get
@@
From the definition of ${\tilde{f}}\left({\mathbf{x}}_{i},{\mathbf{x}}_{i}^{t},\xi_{i}^{t}\right)$ (2), and using (4), we get $\nabla{\tilde{f}}\left({\hat{\mathbf{x}}}_{i}^{t},{\mathbf{x}}_{i}^{t},\xi_{i}^{t}\right)-{\mathbf{z}}_{i}^{t}=\mu_{i}({\hat{{\mathbf{x}}}}_{i}^{t}-{\mathbf{x}}_{i}^{t})$, substituting this gives us:
@@
In $(i)$ and $(ii)$, we have applied the zeroth order convexity condition of $h+11_{\mathcal{X}}$, property of ${\mathbf{W}}$ being doubly stochastic, and $W_{i,j}>0$ for all $i,j\in{\mathcal{V}}$ (A7).
@@
Dividing (31) by $\alpha$ and adding in (30) we get
@@
$\leq\frac{1}{\alpha n}\sum_{i=1}^{n}\left(h\left({\mathbf{x}}_{i}^{t}\right)+11_{{\mathcal{X}}}\left({\mathbf{x}}_{j}^{t}\right)\right)+\frac{1}{n}\sum_{j=1}^{n}\langle\hat{{\mathbf{w}}}_{j}^{t},{\hat{{\mathbf{x}}}}_{i}^{t}-{\mathbf{x}}_{i}^{t}\rangle+\frac{-1}{n}\sum_{i=1}^{n}\mu\left\|{\hat{{\mathbf{x}}}}_{i}^{t}-{\mathbf{x}}_{i}^{t}\right\|^{2}.$
@@
Further applying Cauchy Schwartz inequality and peter-paul’s inequality for $\gamma_{1}>0$ we get,
@@
Taking the Expectation on both sides and summing over $1\leq t\leq T$
@@
Further from the zeroth order convexity condition of $h({\mathbf{x}})$ we have $h({\bar{{\mathbf{x}}}}^{T+1})\leq\frac{1}{n}\sum_{i=1}^{n}h({\mathbf{x}}_{i}^{T+1})$ and from the initialization of D-MSSCA we have ${\mathbf{x}}_{i}^{1}={\bar{{\mathbf{x}}}}^{1}$ for all $i\in{\mathcal{V}}$, using which we get the desired result.